At the local fast food restaurant, 5 cokes and 7 orders of french fries cost $7.25, while 4 cokes and 5 orders od french fries cost $5.50. Then what are the prices of a single coke and a single order of fries?
Answer by nerdybill(7384)

At the local fast food restaurant, 5 cokes and 7 orders of french fries cost $7.25, while 4 cokes and 5 orders od french fries cost $5.50. Then what are the prices of a single coke and a single order of fries?
.
Let c = cost of single coke
and f = cost of single french fry
then our system of equations is
5c + 7f = 7.25
4c + 5f = 5.50
.
multiply top eq by 4 and bottom eq by -5 then add together:
20c + 28f = 29
-20c - 25f = -27.5
-------------------
3f = 1.50
f= .50 (cost of single order of french fry)
.
cost of single coke:
plug above into:
4c + 5f = 5.50
4c + 5(.50) = 5.50
4c + 2.50 = 5.50
4c = 3.00
c = .75